# Consistent Validation for Predictive Methods in Spatial Settings

### v1

@@ 6 Experiments
We next provide experiments showing that the holdout and 1NN struggle in certain test-risk prediction tasks while SNN remains competitive across tasks. To be able to access ground truth, we mainly focus on simulations. But we also show the results of a real-data air temperature modelling task. In all experiments, we run SNN with the probability threshold $\delta=0.1$ and the loss upper bound $\Delta=1$.
6.1 Risk Estimation on Synthetic Data
@@
We set up two synthetic experiments: a grid prediction task and a point prediction task. Based on our analyses above, we expect the holdout to struggle with the former and 1NN to struggle with the latter; our experiments confirm this intuition. See Section C.1.1 for full experiment details.
@@
Figure 2: Error for test-risk estimation in the grid task across methods (holdout in blue, 1NN in orange, our SNN in green); lower values correspond to better performance. The vertical axis shows the absolute difference between the estimated test risk and empirical test risk. Each box plot shows the median, inter-quartile range, and outliers based on 100 synthetic datasets. The horizontal axis tracks increasing validation set sizes.
@@
Validation and Test Data. In both experiments, we vary $N^{\textup{{val}}}\in\{250,500,1000,2000,4000,8000\}$. We use a truncated squared loss: $\ell(a,b)=\max((a-b)^{2},1.0)$. For the grid task, the test sites comprise a $50\times 50$ grid of equally spaced points in $[-0.5,0.5]^{2}$ (orange points in Fig. 1). We generate the validation sites via a sequential process that leads to clustering (blue points in Fig. 1). For the point task, there is a single test site at $(0,0)$. Validation sites are i.i.d. uniform in $[-0.5,0.5]^{2}$. For both tasks, we generate covariates and responses conditional on the sites:
$\!Y_{i}^{\smash{j}}$ $=f(\chi^{1}(S^{\smash{j}}_{i}),\chi^{2}(S^{\smash{j}}_{i}))+\eta(S^{j}_{i})+\epsilon_{i}^{\smash{j}},\,\epsilon^{j}_{i}\sim N(0,\sigma^{2})$
$\eta$ $\sim\mathcal{GP}(0,k_{1}),\,\chi^{1},\chi^{2}\stackrel{{\scriptstyle\textup{{iid}}}}{{\sim}}\mathcal{GP}(0,k_{2}),\,f\sim\mathcal{GP}(0,k_{3}),$
where $\mathcal{GP}(\mu,k)$ denotes a Gaussian process (GP) with mean function $\mu$ and kernel $k$. We describe our kernel and parameter choices in Section C.1.1. We plot examples of the generated data in Figs. 6 and 8.
@@
Figure 3: Error for test-risk estimation in the point task across methods. Axes and box plot details are as in Fig. 2.
Ground Truth. We treat the empirical test risk $\hat{R}^{\chi}_{Q^{\textup{{test}}}}(h)$ as ground truth, with
$\hat{R}^{\chi}_{Q^{\textup{{test}}}}(h):=(1/M^{\textup{{test}}})\sum\nolimits_{m=1}^{M^{\textup{{test}}}}\ell(Y^{\textup{{test}}}_{m},h^{\chi}(S^{\textup{{test}}}_{m}))$ (15)
an unbiased Monte Carlo estimate of the test risk. To compute the empirical test risk, we take $2500$ samples of the covariates and responses at the test sites. [^7] For the point task, we generate 2500 $Y^{\textup{{test}}}_{i}$ values at the origin. By Hoeffding’s inequality, the empirical test risk is very close to the true test risk; namely, with probability $0.95$, $|\hat{R}^{\chi}_{Q^{\textup{{test}}}}(h)-R^{\chi}_{Q^{\textup{{test}}}}(h)|<0.028$. We expect even closer agreement in practice. Figs. 2 and 3 show the absolute difference between each estimator and the empirical test risk across 100 seeds.
@@
Results. We expect our SNN estimator to be consistent in all tasks. In the grid task, we expect the variance of 1NN to be low since there are many, dense test points. And we expect the bias of the holdout to be high since the test and validation points have noticeably different spatial arrangement. Our results in the grid task (Fig. 2) agree with our intuitions; the errors of 1NN and our SNN decrease much more rapidly across the values of $N^{\textup{{val}}}$ than the holdout.
@@
In the point task, we expect the high variance of 1NN to be an issue with substantial probability given that there is only one test point. Fig. 3 agrees with our intuition; the error of our SNN estimator decreases much more rapidly across the values of $N^{\textup{{val}}}$ than 1NN. In this case, we find that the holdout errors decrease rapidly as well. We also plot the (signed) relative difference of each estimator to the empirical test risk in Figs. 7 and 9.
@@
In Tables 2 and 3, we show the number $k$ of nearest neighbors selected by SNN for each of the two tasks. In the grid task, the optimal $k$ was at most 2 in all cases we considered. In the point task, we found the optimal $k$ noticeably increased with $N^{\textup{{val}}}$, though it always remained over an order of magnitude smaller than $N^{\textup{{val}}}$.
6.2 Model Selection on Synthetic Data
We next see that SNN and 1NN are able to select the model with lower test risk in a model selection task, but the holdout systematically picks the wrong model. We repeat the model selection problem 100 times. In each repetition, we have $N^{\text{train}}=100$ and a max $N^{\textup{{val}}}=75$. In our analysis, we will consider validation subsets of size $N^{\textup{{val}}}\in\{5\ell\}_{\ell=1}^{15}$. We generate independent test, validation, and training data as follows; see Fig. 4.
$U_{i}^{\smash{j}}$ $\sim\mathcal{U}([0,1]),\quad S_{i}^{\smash{j}}=\sqrt{U_{i}^{\smash{j}}},\quad j\in\{\text{train},\text{val}\}$
$S_{m}^{\smash{\text{test}}}$ $=m/20,0\leq m\leq 20,\quad\epsilon_{\smash{i}}^{\smash{j}}\sim\mathcal{U}([0,0.1]),$
$Y_{\smash{i}}^{j}$ $=|S_{\smash{i}}^{j}-0.5|+\epsilon_{\smash{i}}^{j}\quad j\in\{\text{train},\text{val},\text{test}\},$ (16)
We compare two predictive methods: $h_{0}(S)=0.25$ and $h_{1}(S)=\beta_{1}^{\top}S+\beta_{0}$, with $(\beta_{1},\beta_{0})$ fit by minimizing the mean absolute residual on the training data. Fig. 4 shows the data and predictions of both models (as functions of space). We use the loss function $\ell(a,b)=|a-b|$, which is bounded for this problem because both the hypotheses and the response variable are bounded on $[0,1]$.
Across all seeds, $h_{0}$ has the lower empirical test risk; $h_{1}$ makes large errors on the test points near 00 because most of the training data is in $[0.5,1]$. Since most of the validation data also clusters near 1, we expect the holdout to struggle due to bias. Our arguments in Section 5.3 lead us to expect both SNN and 1NN should perform well on this task when given sufficient validation data.
We say an estimator of the risk, $\hat{R}$, selects $h_{0}$ if $\hat{R}(h_{0})<\hat{R}(h_{1})$. We plot the percentage of times each method (correctly) selects $h_{0}$ as a function of the number of validation points in Fig. 5. When the validation set is small, all estimators select the model with lowest test risk ($h_{0}$) less than half the time. For the nearest neighbor methods, we expect that when there are few or no spatial locations less than $1$, weighting cannot fix the estimate. However, when the number of validation points is large, the nearest neighbor methods consistently (correctly) select $h_{0}$. By contrast, the holdout consistently (incorrectly) selects $h_{1}$, even though $h_{1}$ has higher test risk. See Section C.2 for full experiment details.
6.3 Air Temperature Prediction
Although we do not have access to ground truth in real data, we next show a real-data experiment where the holdout, 1NN, and our SNN give very different estimates of test risk and can differ in model selection.
We consider predicting monthly average air surface temperature based on remote sensing observations of monthly average land surface temperature at day and night from MODIS-Aqua (Wan et al., 2021). This task is inspired by Hooker et al. (2018), who used two weighted regression methods to construct a global map of air temperature.
Our test task is prediction at the 5 largest urban areas in the United States (New York City, Los Angeles, Chicago, Miami, and Houston). Loss is absolute error in degrees Celsius. For predictive methods, we consider geographically weighted linear regression (GWR) and a baseline GP spatial interpolation model fit only to weather station data.
Table 1 shows a large discrepancy in test-risk estimates across estimators. We see that 1NN chooses a different predictive method (spatial GP) than holdout or our SNN does. Given the results in the rest of this paper, we would favor decisions made with SNN rather than the other two estimators. See Section C.3 for full details and also a grid-prediction task where all estimators are in agreement.
@@ 6.3 Air Temperature Prediction
Table 1: Test-risk estimates for the 5-metros task. Rows correspond to predictive methods and columns to estimators. We report two standard error intervals for holdout. For each estimator, we bold the predictive method with lower estimated test risk.
Holdout 1NN SNN
GWR $\mathbf{0.84\pm 0.03}$ $0.60$ $\mathbf{0.67}$
Spatial GP $0.96\pm 0.04$ $\mathbf{0.40}$ $0.75$
7 Discussion
We proposed an adaptive nearest neighbors method for predictive validation. We showed that past validation methods can be inconsistent under infill asymptotics but our method is consistent. A number of open questions remain for spatial validation, such as providing checks when an infill assumption is inappropriate or checks on extrapolation quality.
@@ Acknowledgements
Appendix A Extended Related Work
@@
A.1 Validation versus Cross-Validation
@@
While we expect our work to have implications for cross validation (CV), we focus on validation here since (1) we do not assume training data exists or is easy to change and (2) CV presents additional subtle challenges. The holdout is broadly applicable to any predictive method (whether data-driven, physical, or a combination thereof). However, in the cases when data is scarce and a data-driven predictive method is used, cross-validation is commonly believed to make more efficient use of the data. CV is widely used in spatial analyses; among many examples are Wang et al. (2022); Valavi et al. (2021); Kianian et al. (2021). However, the interpretation of the error estimates given by cross-validation is subtle even in the classical i.i.d. setting (Bates et al., 2023). We focus on the validation setting in this work due to its broad applicability and clear interpretation, though extensions to cross-validation, as well as a theoretical understanding of how the resulting error estimate relate to predictive performance is a promising direction for future work.
A.2 Covariate Shift
@@
In covariate shift, it is generally assumed that $S^{\textup{{val}}}_{n}\stackrel{{\scriptstyle\textup{{iid}}}}{{\sim}}P^{\text{val}}$, $S^{\textup{{test}}}_{m}\stackrel{{\scriptstyle\textup{{iid}}}}{{\sim}}Q$ and $U^{\ell}_{i}|S^{\ell}_{i}\stackrel{{\scriptstyle\textup{{iid}}}}{{\sim}}\mu$ for $\ell\in\{\text{val},\text{test}\}$. Moreover, it is typically assumed the density ratio $\frac{dQ}{dP}$ exists and is bounded, although there is recent work relaxing this assumption in the context of non-parametric regression (Kpotufe & Martinet, 2021; Pathak et al., 2022). Mean estimation seeks to estimate $\mathbb{E}[U^{\text{test}}_{1}]$. Taking $(U^{\text{test}}_{m}|S^{\textup{{test}}}_{m})=\ell(Y^{\textup{{test}}}_{m},h^{\chi}(S^{\textup{{test}}}_{m}))$, this is the same task we consider, but with the (stronger) assumption that covariates are independent and identically distributed with a bounded density ratio between the test and validation distributions. This assumptions is not appropriate for validation with many spatial datasets, with particularly simple examples being when the task of interest requires prediction at a single spatial location, or on a regular grid. Our assumptions are essentially a conditional formulation of mean estimation under covariate shift. Many methods proposed for addressing mean estimation under covariate shift, including all the approaches we describe below, are based on re-weighting validation points, then applying the holdout approach described earlier with these weights.
@@
The kernel mean matching algorithm provides a solution under standard covariate shift assumptions with a bounded density ratio and assuming conditional expectation $e^{\chi}_{h}$ lives in a reproducing kernel Hilbert space (RKHS) and has small norm (Gretton et al., 2009). Yu & Szepesvári (2012) provide a finite sample bound for this method
Appendix B Proofs of Claims

### v2

@@ 6 Experiments
Our theory suggests the holdout exhibits substantial bias in many tasks. And we expect 1NN to exhibit substantial variance in many point prediction tasks. Our experiments confirm these observations. While there exist tasks where either the holdout or 1NN performs similarly to SNN, there are also many tasks where each performs much worse than SNN. Since SNN performs well across all experiments, we prefer SNN when a new task arises.
Ground Truth. In a traditional machine learning prediction task, analysts ask how well a predictive method predicts the observed response at a set of covariates, so the observed response would form the ground truth. Since here we instead judge evaluation methods, we must ask how well the evaluation method estimates test risk (Definition 2.3); that is, the true test risk now forms ground truth. From Assumption 2.4, the true test risk requires an integral over the (unknown) noise $P_{\epsilon}$. Accessing ground-truth responses is often easy; by contrast, it is highly unusual to access even a high-quality approximation of the integral for test risk (much less the exact test risk) in a real task. Therefore we devise a series of workarounds. First, we consider realistic tasks with simulated data. Second, we consider a realistic task with a semi-simulated data set, where we control the noise distribution by constructing it from bootstrapped residuals that arise from real data. Third, we use fully real data to construct a ground truth by considering an unrealistic task. Finally, we consider fully real data and a realistic task by forfeiting access to ground truth.
6.1 Test Risk Estimation on Fully Synthetic Data
@@
We set up two fully synthetic experiments: a grid prediction task and a point prediction task. Based on our analyses above, we expect the holdout to struggle with the former and 1NN to struggle with the latter; our experiments confirm this intuition. See Section E.3 for full experiment details.
@@
Figure 1: Error for test risk estimation in the grid prediction task (left) and point prediction task (right) across methods (holdout in blue, 1NN in orange, our SNN in green); lower values correspond to better performance. The vertical axis shows the absolute difference between the estimated test risk and empirical test risk. Each box plot shows the median, inter-quartile range, and outliers based on 100 synthetic datasets. The horizontal axis tracks increasing validation set sizes. Numbers above the upper box indicate the number of outliers falling above the vertical limit.
@@
Validation Data, Test Data, and Ground Truth. In both experiments, we vary $N^{\textup{{val}}}\in\{250,500,1000,2000,4000,8000\}$. We use a truncated squared loss: $\ell(a,b)=\max((a-b)^{2},1.0)$. For the grid task, the test sites comprise a $50\times 50$ grid of equally spaced points in $[-0.5,0.5]^{2}$ (orange points in Figure 3). We generate the validation sites via a sequential process that leads to clustering (blue points in Figure 3). For the point task, there is a single test site at $(0,0)$. Validation sites are i.i.d. uniform in $[-0.5,0.5]^{2}$. For both tasks, we generate covariates and responses conditional on the sites: $Y_{i}^{\smash{j}}=f(\chi^{1}(S^{\smash{j}}_{i}),\chi^{2}(S^{\smash{j}}_{i}))+\eta(S^{j}_{i})+\epsilon_{i}^{\smash{j}},\,\epsilon^{j}_{i}\smash{\stackrel{{\scriptstyle\textup{{\tiny iid}}}}{{\sim}}}N(0,\sigma^{2})$. We generate $\eta,\chi^{1},\chi^{2},f$ according to independent Gaussian processes (GPs); we describe our kernel and parameter choices in Section E.3.1. We plot examples of the generated data in Figures 4 and 6. We make draws from the data-generating process to form an unbiased Monte Carlo estimate of the test risk, $\smash{\hat{R}_{Q^{\textup{{test}}}}}(h)$, and use this estimate as ground truth; see Section E.3.5 for details.
@@
Results. We expect our SNN estimator to be consistent in all tasks. In the grid task, we expect the variance of 1NN to be low since there are many test points spread across the domain. And we expect the bias of the holdout to be high since the test and validation points have noticeably different spatial arrangement. Our results in the grid task (Figure 1, left) agree with our intuitions; the errors of 1NN and our SNN decrease much more rapidly across the values of $N^{\textup{{val}}}$ than the holdout.
@@
Given the single test point in the point task, we expect the high variance of 1NN to be an issue with substantial probability. Figure 1 (right) agrees with our intuition; the error of our SNN estimator decreases much more rapidly across the values of $N^{\textup{{val}}}$ than 1NN. In this case, we find that the holdout errors decrease rapidly as well. We also plot the (signed) relative difference of each estimator to the empirical test risk in Figures 5 and 7 (Section E.3).
@@
In Tables 2 and 3 (Section E.3), we show $k$, the number of nearest neighbors selected by SNN for each of the two tasks. In the grid task, the $k$ selected was at most 4 in all cases we considered. In the point task, the value of $k$ selected increased with $N^{\textup{{val}}}$, though it always remained over an order of magnitude smaller than $N^{\textup{{val}}}$.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Signed errors in estimating the test risk for (left to right) the bootstrapped experiment with geographically weighted regression; the same task with GP regression; the flat price prediction task; and the wind speed prediction task. The holdout (blue), 1NN (orange), and SNN (green) appear left to right in each plot.
@@
Table 1: Test risk estimates for the 5-metros task. Rows correspond to predictive methods and columns to estimators. We report two standard error intervals for holdout. For each estimator, we bold the predictive method with lower estimated test risk.
GWR Spatial GP
Hold.
$\mathbf{0.83}$
±$\mathbf{\pm 0.03}$
$\mathbf{0.90}$
±$\mathbf{\pm 0.04}$
1NN $0.61$ $\mathbf{0.44}$
SNN $\mathbf{0.53}$ $0.61$
6.2 Air Temperature Task with Bootstrapped Residuals
We next consider a real task on a semi-synthetic dataset. We find that 1NN performs poorly; while the holdout performs best, SNN performs well. Full details can be found in Section E.4.
Data and Ground Truth. Our test task is prediction of monthly average air temperature in January 2023 at the 5 largest urban areas in the United States (New York City, Los Angeles, Chicago, Miami, and Houston), based on available weather station data in the same month (Menne et al., 2018). Loss is truncated absolute error (in °C). To access ground truth test risk, we create a partially synthetic response variable. We first fit Gaussian process regression (GPR) to all the available weather station data. We build 100 synthetic datasets by calculating the residuals of the posterior mean of this model, sampling a residual value for each weather station and point we want to predict at and adding these to the mean prediction of the model. Because we then have access to samples from the distribution of response values at the test sites under this data-generating process, we can obtain a (accurate estimate of) ground truth test risk of a predictive method on this partially synthetic response (Section E.4.7).
Spatial Predictive Methods. We train two predictive methods on this data: GPR and a geographically weighted regression (GWR) based on MODIS-Aqua (Wan et al., 2021) land surface temperature measurements, inspired by Hooker et al. (2018). We use 50% of the weather station locations for training the predictive methods and the remaining 50% for validation (3211 observations in each).
Results. The error in estimating the predictive performance of both methods is shown in Figure 2 (far and mid left) for 100 different datasets with different samples of the residuals (but the same training and validation split). Given the point prediction task, we expect 1NN (orange, middle) to have a high variance; the figure confirms our intuition. The estimates given by SNN (green, right) and the holdout (blue, left) are much closer to the ground truth. In this case, the holdout has a small bias, and its variance is substantially lower than SNN. So, in this case, the holdout typically returns slightly better estimates of the error than SNN. If many more weather stations were used for validation, we expect that the SNN would eventually outperform the holdout estimate.
6.3 Property Sales in England and Wales
Here and in Section 6.4, we define somewhat unrealistic tasks to access ground truth on fully real data. Here we find that 1NN and SNN perform similarly well while the holdout exhibits a large bias. Section E.5 contains additional details and figures for this experiment.
Data and Ground Truth. We consider prediction of the price of a flat in England and Wales based on location, loosely following Hensman et al. (2013) but using data from 2023 (HM Land Registry, 2023). We make 100 datasets by sampling a training dataset of 40,000 points from flat sales outside London, a test set consisting of 1,000 flat sales within London, and a validation set consisting of the remaining sales in 2023 (31,484 outside London, 21,179 in London). The loss is truncated mean absolute error. In Section E.5.3, we justify how we can form a high-quality estimate of ground-truth test risk by: assuming a form of independence, using a bounded loss, and applying Hoeffding’s inequality.
Spatial Predictive Method. We fit a Gaussian process regression model with variational inference as in Hensman et al. (2013) with minor modifications; we use a sum of two Matérn 3/2 kernels. We use the non-stochastic version of variational inference in GPR (Titsias, 2009) to avoid known difficulties with tuning hyperparameters in the stochastic version (Ober et al., 2024), and 2000 inducing points.
Results. Because the model is trained using data only outside London, we expect it to make larger errors predicting flat sales in London than outside London. As a result, we expect the holdout to have a large bias. We expect both 1NN and SNN to perform similarly: since the test sites are sampled randomly within London, we expect them to have different nearest neighbors and so variance of 1NN should be reasonably small. The mean absolute error of each method, relative to the estimate of the ground truth, is shown in Figure 2 (mid right). As expected, the holdout substantially underestimates the test risk, while the other estimators perform reasonably well.
6.4 Wind Speed Prediction
In this next experiment with an unrealistic task but fully real data, we find that both the holdout and 1NN perform poorly while SNN performs well. Full details are in Section E.6.
Data and ground truth. Our test task is predicting the average wind speed on a typical day in January at Chicago O’Hare airport, from daily historical weather station data (Menne et al., 2012). There are 775 historical weather station observations at the Chicago O’Hare site in January months. We split the remaining weather stations (at a station level) into a training and validation set consisting of 962 training stations and 241 validation stations. Each station has a different number of observations depending on how complete the historical record of average daily wind speed data is at this site. Typically the training set consists of on the order of $580{,}000$ measurements and the validation set around $126{,}000$ measurements. We perturb the location labels of each measurement at each weather station by a tiny amount to avoid ties when running nearest neighbors. The test set contains all of January rather than just a particular date so that we can form a high-quality estimate of ground truth. To form ground truth, we also assume that average wind speed decorrelates in time quickly. The loss is truncated (root) mean square error. See Section E.6.3 for more details.
Spatial Predictive Method. We use LightGBM (Ke et al., 2017) to make predictions.
Results. We expect 1NN to have a high variance because the task is point prediction. Figure 2 (right) confirms this. The holdout has large bias for this task while SNN exhibits low bias and low variance.
6.5 Air Temperature Prediction with Real Response
We finally consider a case with real data and a real task. Although we cannot access ground truth, we show that the holdout, 1NN, and our SNN give very different estimates of test risk and can differ in model selection. Given all the previous results, we suggest using SNN. See Section E.4 for full details and also a grid-prediction task where all estimators are in agreement.
Data and Models Fit. The data, test task and models considered are the same as in Section 6.2, but the actual response values are used for training and validation.
Results. Table 1 shows a large discrepancy in test risk estimates across estimators. We see that 1NN chooses a different predictive method (spatial GP) than the holdout or our SNN does.
@@ Acknowledgements
Appendix
[…]
Appendix A Broader Impact Statement
This paper identifies potential failure modes of existing validation techniques in a spatial setting and suggests practical improvements, with supporting theory and empirics. We hope that our work can play a part in improving the reliability of prediction assessment and thereby help improve the credibility of scientific analyses.
Appendix B Extended Related Work
@@
B.1 Validation versus Cross-Validation
@@
While we expect our work to have implications for cross validation (CV), we focus on validation here since (1) we do not assume training data exists or is easy to change (for example in the case of a large physical model) and (2) CV presents additional subtle challenges. The holdout is broadly applicable to any predictive method (whether data-driven, physical, or a combination thereof). However, in the cases when data is scarce and a data-driven predictive method is used, cross-validation is commonly believed to make more efficient use of the data. CV is widely used in spatial analyses; among many examples are Wang et al. [2022], Valavi et al. [2021], Kianian et al. [2021]. However, the interpretation of the error estimates given by cross-validation is subtle even in the classical i.i.d. setting [Bates et al., 2023]. We focus on the validation setting in this work due to its broad applicability and clear interpretation, though extensions to cross-validation, as well as a theoretical understanding of how the resulting error estimate relate to predictive performance is a promising direction for future work.
B.2 Covariate Shift
@@
In covariate shift, it is generally assumed that $S^{\textup{{val}}}_{n}\stackrel{{\scriptstyle\textup{{\tiny iid}}}}{{\sim}}P^{\text{val}}$, $S^{\textup{{test}}}_{m}\stackrel{{\scriptstyle\textup{{\tiny iid}}}}{{\sim}}Q$ and $U^{\ell}_{i}|S^{\ell}_{i}\stackrel{{\scriptstyle\textup{{\tiny iid}}}}{{\sim}}\mu$ for $\ell\in\{\text{val},\text{test}\}$. Moreover, it is typically assumed the density ratio $\frac{dQ}{dP}$ exists and is bounded, although there is recent work relaxing this assumption in the context of non-parametric regression [Kpotufe and Martinet, 2021, Pathak et al., 2022]. Mean estimation seeks to estimate $\mathbb{E}[U^{\text{test}}_{1}]$. Taking $(U^{\text{test}}_{m}|S^{\textup{{test}}}_{m})=\ell(Y^{\textup{{test}}}_{m},h^{\chi}(S^{\textup{{test}}}_{m}))$, this is the same task we consider, but with the (stronger) assumption that covariates are independent and identically distributed with a bounded density ratio between the test and validation distributions. This assumptions is not appropriate for validation with many spatial datasets, with particularly simple examples being when the task of interest requires prediction at a single spatial location, or on a regular grid. Our assumptions are essentially a conditional formulation of mean estimation under covariate shift. Many methods proposed for addressing mean estimation under covariate shift, including all the approaches we describe below, are based on re-weighting validation points, then applying the holdout approach described earlier with these weights.
@@
The kernel mean matching algorithm provides a solution under standard covariate shift assumptions with a bounded density ratio and assuming the average loss as a function of space $e_{h}$ lives in a reproducing kernel Hilbert space (RKHS) and has small norm [Gretton et al., 2009]. Yu and Szepesvári [2012] provide a finite sample bound for this method
Appendix C Error Estimation for Model Selection
(1) Fixed predictive methods. So far we have focused on the setting where predictive methods are fixed in advance. As we accrue validation data, the consistency of SNN ensures it will eventually choose the predictive method with lowest test risk. Without consistency, holdout and 1NN cannot be trusted to choose the best predictive method.
We are not able to formally characterize how SNN performs in model selection. But we provide rigorous results on rates of convergence of SNN that give suggestive guidance. In particular, (a) we first consider the case of a grid prediction task, where test sites are arranged on a grid. In this case, we are able to give a rate of convergence for both $1$NN and SNN (Corollaries D.16 and D.17). If the validation data are i.i.d. our rate is faster than the minimax optimal convergence rate for predicting a Lipschitz function in the presence of additive, homoskedastic noise. Our result shows the same rate of convergence as Portier et al. [2023, Prop. 4], who considered test data that was i.i.d. instead of on a grid. Our result is suggestive that both nearest neighbor methods may perform well at model selection for validating maps (grid prediction).
(b) Second, we provide finite-sample bounds and asymptotic characterizations for general (non-grid) $Q^{\textup{{test}}}$ (Corollary D.14). In this case, when validation data is i.i.d., up to logarithmic factors, we show that SNN converges at the optimal rate of convergence for Lipschitz functions (Corollary D.18). Since in this case, the SNN convergence rate is not strictly faster than the training convergence rate, it might be difficult to select between statistical methods that converge at the optimal regression rate. But we would still expect to be able to select between (i) a statistical method that converges at the optimal rate and (ii) one that does not (for example a misspecified parametric model). Since the holdout and 1NN may not even be consistent, we could not rely on them to select the better model even in this latter, easier case.
Appendix D Proofs of Claims